Find the value of k so that the line 3x-5+k=0 may touch the parabola y2=24x?

Grade:

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear LINTA

equation of parabola   y2=4*6*x

                       a =6

equation of given equation  3x-5y+k=0

                                                or   y = 3x/5  + k/5

                                  m = 3/5    and  c= k/5

condition for tangency   C =a/m

                                           k/5  = 6/(3/5)

                                           k = 50

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Nehal Wani
14 years ago
y2=4*6*x a =6 equation of given equation 3x-5y+k=0 or y = 3x/5 + k/5 m = 3/5 and c= k/5 condition for tangency C =a/m k/5 = 6/(3/5) k = 50